# Concurrent Composition for Differentially Private Continual Mechanisms

### v3

@@ 8. Reduction to Interactive Randomized Response
Lyu [19] shows that for every $(\epsilon,\delta)$-DP IM $\mathcal{M}$ and every pair of neighboring initial states $s_{0},s_{1}$ for $\mathcal{M}$, there exists an IPM $\mathcal{P}$ such that for each $b\in\{0,1\}$, the mechanisms $\mathcal{M}(s_{b})$ and $\mathcal{P}\!\circ^{*}\!\mathrm{RR}_{\epsilon,\delta}(b)$ are equivalent. Meaning that, $\mathcal{P}$ can simulate $\mathcal{M}(s_{b})$ given access only to the output of a randomized response mechanism $\mathrm{RR}_{\epsilon,\delta}(b)$. In this section, inspired by this work, we construct a new IPM $\mathcal{P}$ that, rather than simulating $\mathcal{M}$ based on the outcome of a single instance of $\mathrm{RR}_{\epsilon,\delta}$, operates on the output of $\mathrm{RR}_{0,\delta}$ and only when necessary uses the output of $\mathrm{RR}_{\epsilon,0}$. In other words, instead of interacting with the NIM $\mathrm{RR}_{\epsilon,\delta}$, $\mathcal{P}$ interacts with the IM $\mathrm{IRR}_{\epsilon,\delta}$, and under certain conditions, it queries $\mathrm{IRR}_{\epsilon,\delta}$ at most once.
@@
Lemma (Restatement of Lemma 5.6).
@@
Fixing an upper bound $T$ on the number of queries, Lyu [19] reduces the construction of $\mathcal{P}$ to defining a sequence of functions $\phi_{t}^{0}$ and $\phi_{t}^{1}$ for $t\in[T]$ that satisfy three conditions. They then inductively construct $\phi_{t}^{b}$ and verify that it satisfies the required conditions. The definition of these conditions is based on a sequence of control function $\mathrm{L}_{t,T}^{0}$ and $\mathrm{L}_{t,T}^{1}$, which have backward recursive definitions for $t$ from $T$ to $1$.
@@
In Section 8.2, we summarize the reduction introduced in [19], with the control functions $\mathrm{L}_{t,T}^{b}$ replaced by $\mathrm{L}_{t}^{b}$. We also modify this reduction by adding a new condition to the original three that the functions $\phi_{t}$ must satisfy. This additional condition is used to limit the number of interactions between $\mathcal{P}$ and $\mathrm{IRR}_{\epsilon,\delta}(b)$. Finally, in Section 8.3, we provide a new construction of $\phi_{t}$ and show that it satisfies all four conditions.
@@
Lyu [19] introduces the following control function:
@@
Lyu [19] shows the following two claims about $\mathrm{Lower}_{t,T}^{b}$:
Claim 8.5 ([19]).
For each $b\in\{0,1\}$, $T\in\mathbb{N}$, and $q_{1}\in Q$ $\mathrm{Lower}_{1,T}^{b}(q_{1})\leq\delta$.
Claim 8.6 ([19]).
For each $b\in\{0,1\}$, $T\in\mathbb{N}$, $t\in[T]$, [^7] In the original claim in [19], $t$ is assumed to be at most T1T-1, which is a typo. and transcript $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{t-1}\times Q$, we have
@@
For each $b\in\{0,1\}$, and $T\in\mathbb{N}$, $\mathrm{L}_{0,T}^{b}()\leq\delta$.
In some part of the proof, Lyu [19] wrongly refers to Claim 8.6, while the following lemma is required. For completeness, we prove this lemma and use it in the next section.
@@
This section introduces our main new technical contribution. To remove the assumption of having an upper bound $T$ on the number of queries, we introduce $\mathrm{L}_{t}^{b}$ as limit of $\mathrm{L}_{t,T}^{b}$ and show that it satisfies all properties needed by the proof of [19]. Thus, differently from [19], we will use $\mathrm{L}_{t}^{0}$ and $\mathrm{L}_{t}^{1}$ to define the function $\phi_{t}$.
@@
To ensure $\mathrm{L}_{t}^{b}$ is well-defined, we need to show that the limit of $\mathrm{L}_{t,T}^{b}$ exists. We will upper bound the values of $\mathrm{L}_{t,T}^{b}$ by $1$ and show that this function is non-decreasing in $T$. As a result, for each each $t\in\mathbb{N}\cup\{0\}$ and $(q_{1},\dots,a_{t})\in(Q\times A)^{t}$, the limit $\lim_{T\geq t,T\to\infty}\mathrm{L}_{t,T}^{b}(q_{1},\dots,a_{t})$ exists. We need the following lemma throughout the section:
@@
$=\sum\limits_{a_{T}\in A}\sup_{q_{T+1}\in Q}\max\{0,\sum\limits_{a_{T+1}\in A}\mu_{T+1}^{b}(q_{1},\dots,,a_{T+1})-e^{\epsilon}\mu_{T+1}^{1-b}(q_{1},\dots,a_{T+1})\}$
@@
In this section, we present the properties of $\mathrm{L}_{t}^{b}$, required in the next sections. The definition of $\mathrm{L}_{t}^{b}$ combined with Corollary 8.7, Lemma 8.8, and Lemma 8.12 give the following corollaries, respectively:
@@
For each $b\in\{0,1\}$, $\mathrm{L}_{0}^{b}()\leq\delta$.
@@
To define $\mathcal{P}$ and simulate $\mathcal{M}$, Lyu [19] constructs four families of probability mass functions (PMFs), corresponding to each outcome of $\mathrm{RR}_{\epsilon,\delta}$. Intuitively, each family describes an IM. Specifically, each family describes an IM and consists of PMFs over the answer set $A$ for every $t\in[T]$ and every history $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{t}\times Q$. These PMFs represent the probability of returning each answer in response to $q_{t}$, conditioned on $q_{1},\dots,a_{t-1}$ being the history of queries and answers. Similarly, we define $\mathcal{P}$ based on four families of PMFs: $\mathcal{F}_{(\bot,0)}$, $\mathcal{F}_{(\bot,1)}$, $\mathcal{F}_{(\top,0)}$, and $\mathcal{F}_{(\top,1)}$, corresponding to the four possible output sequences of $\mathrm{IRR}_{\epsilon,\delta}$. We note that in our construction, each family includes a PMF over $A$ for every $t\in\mathbb{N}$ and every history $(q_{1},\dots,a_{t-1},q_{t})$. Moreover, the definitions of these PMFs are not the same as the ones in [19].
@@
In [19], the post-processing mechanism initially interacts with $\mathrm{RR}_{\epsilon,\delta}$ and stores its output $r=(\tau,c)\in\{\top,\bot\}\times\{0,1\}$. It then answers the adversary’s queries by itself using the family of PMFs corresponding to $r$. Our high-level idea is to construct the families $\mathcal{F}_{(\bot,0)}$, $\mathcal{F}_{(\bot,1)}$, $\mathcal{F}_{(\top,0)}$, and $\mathcal{F}_{(\top,1)}$ in such a way that, under certain conditions, knowledge of $\tau$ alone suffices to choose the appropriate PMF. Recall that the IPM $\mathcal{P}$ interacts with $\mathrm{IRR}_{\epsilon,\delta}$ instead of $\mathrm{RR}_{\epsilon,\delta}$. We design $\mathcal{P}$ to initially interact with $\mathrm{IRR}_{\epsilon,\delta}$ and store its response. $\mathcal{P}$ then interacts with $\mathrm{IRR}_{\epsilon,\delta}$ for the second time when it cannot choose a PMF using only the first response.
@@
To generate a response $a$ to query $q$, $\mathcal{P}$ proceeds as follows: If $q$ is the first query (i.e., $r=()$), then $\mathcal{P}$ initially interacts with $\mathrm{IRR}_{\epsilon,\delta}$ and appends its response to $r$. After this, whether $q$ is the first query or not, when the following condition does not hold for the first time, $\mathcal{P}$ interacts with $\mathrm{IRR}_{\epsilon,\delta}$ for the second time and appends its answer to $r$: the PMFs corresponding to $(h,q)$ in $\mathcal{F}_{(\top,0)}$ and $\mathcal{F}_{(\top,1)}$ are identical, and the PMFs corresponding to $(h,q)$ in $\mathcal{F}_{(\bot,0)}$ and $\mathcal{F}_{(\bot,1)}$ are identical. After that, $\mathcal{P}$ never interacts with $\mathrm{IRR}_{\epsilon,\delta}$ again.
@@
Using the control function $\mathrm{Lower}_{t,T}$, Lyu [19] introduces three conditions and constructs a sequence of functions $\phi_{t}:(Q\times A)^{t}\to[0,1]^{2}$ for $t\in\{1,\dots,T\}$ that satisfy these conditions. The families of PMFs are defined then based on these functions. We remove the dependence on the bound $T$ by replacing $\mathrm{Lower}_{t,T}$ with $\mathrm{L}_{t}$ in the conditions, and extend the construction of $\phi_{t}$ to all $t\in\mathbb{N}$. Furthermore, we introduce an additional condition to those given in [19], which we will later use to bound the number of times $\mathcal{P}$ interacts with $\mathrm{IRR}_{\epsilon,\delta}$.
@@
Similar to the previous section, $\mathcal{F}_{(0,\top)}$ and $\mathcal{F}_{(1,\top)}$ are defined based on a sequence of functions $\psi_{t}:(Q\times A)^{t}\to[0,1]^{2}$ for each $t\in\mathbb{N}\cup\{0\}$. For $t=0$, $\psi_{t}()=(1,1)$. For $t\in\mathbb{N}$, we define $\psi_{t}$ as follows:
@@ 8.3. Construction of $\phi_{t}$
For each $q_{1},\dots,a_{t-1},q_{t}$, Lyu [19] initializes the values of $\phi_{t}(q_{1},\dots,a_{t-1},q_{t},\cdot)$ with the ones from $\mathrm{L}_{t,T}(q_{1},\dots,a_{t-1},q_{t},\cdot)$. (Removing the upper bound assumption, they are initialized by the values of $\mathrm{L}_{t}(q_{1},\dots,a_{t-1},q_{t},\cdot)$.) Then, they iterate over all possible answers $a_{t}\in A$ in a fixed order and gradually increase the value of $\phi_{t}(q_{1},\dots,a_{t})$ to satisfy conditions (II) and (III), while ensuring that condition (I) remains valid. This iterative procedure requires $A$ to be finite; otherwise, the construction algorithm does not terminate. Moreover, the fixed iteration order can lead to an imbalance in which earlier values grow significantly larger than those that come later in the order. In our new construction, we control and balance this increase.
@@
Lemma 8.25 (Transfinite Recursion [22]).
@@ 8.3.4. Satisfying Conditions (I) to (IV)
Our objective was to construct $\phi_{t}^{b}$ so that it fulfills conditions (I) to (IV). Combination of Constraints ($\phi$.1) and ($\xi$.1) gives condition (I). Moreover, Constraint ($\phi$.3) directly imply condition (II). The proof of the following claim is identical to a proof in [19]:
@@
which is a contradiction when $\epsilon>0$, which implies that condition (III) holds. ∎
@@
for all $(a_{1},\dots,a_{t})\in A^{t}$. Moreover, by Lemma 8.10 and $\star$8.3.4, $\mu_{t-1}^{0}(q_{1},\dots,a_{t-1})=\mu_{t-1}^{1}(q_{1},\dots,a_{t-1})$ for all $(a_{1},\dots,a_{t})\in A^{t}$. Thus, by Remark 8.21 and condition (IV),
@@
Therefore, by equations ($\star$8.3.4), ($\star\star$8.3.4) and ($\star\star\star$8.3.4), Constraints ($\phi$.1), ($\phi$.2), and ($\phi$.3) are symmetric for $x_{0}$ and $x_{1}$. Hence, there exists a maximal feasible solution $(z^{*},z^{*})$ satisfying all the constraints. By definition of $\phi_{t}(q_{1},\dots,a_{t})$, we have $\phi_{t}(q_{1},\dots,a_{t})=(z^{*},z^{*})$. Also, by Constraint ($\phi$.1), $z^{*}$ is at least $\min\{\mu_{t}^{0}(q_{1},\dots,a_{t}),\mathrm{L}_{t}^{0}(q_{1},\dots,a_{t})+\mathrm{L}_{t}^{1}(q_{1},\dots,a_{t})\}$.

### v4

@@ 10. Reduction to Interactive Randomized Response
Lyu (lyu2022composition, ) shows that for every $(\epsilon,\delta)$-DP IM $\mathcal{M}$ and every pair of neighboring initial states $s_{0},s_{1}$ for $\mathcal{M}$, there exists an IPM $\mathcal{P}$ such that for each $b\in\{0,1\}$, the mechanisms $\mathcal{M}(s_{b})$ and $\mathcal{P}\!\circ^{*}\!\mathrm{RR}_{\epsilon,\delta}(b)$ are equivalent. This means that $\mathcal{P}$ can simulate $\mathcal{M}(s_{b})$ given access only to the output of a randomized response mechanism $\mathrm{RR}_{\epsilon,\delta}(b)$. In this section, inspired by this work, we construct a new IPM $\mathcal{P}$ that, rather than simulating $\mathcal{M}$ based on the outcome of a single instance of $\mathrm{RR}_{\epsilon,\delta}$, operates on the output of $\mathrm{RR}_{0,\delta}$ and only when necessary uses the output of $\mathrm{RR}_{\epsilon,0}$. In other words, instead of interacting with the NIM $\mathrm{RR}_{\epsilon,\delta}$, $\mathcal{P}$ interacts with the IM $\mathrm{IRR}_{\epsilon,\delta}$, and under certain conditions, it queries $\mathrm{IRR}_{\epsilon,\delta}$ at most once.
@@
Lemma 0 (Restatement of Lemma 5.6).
@@
Fixing an upper bound $T$ on the number of queries, Lyu (lyu2022composition, ) reduces the construction of $\mathcal{P}$ to defining a sequence of functions $\phi_{t}^{0}$ and $\phi_{t}^{1}$ for $t\in[T]$ that satisfy three conditions. They then inductively construct $\phi_{t}^{b}$ and verify that it satisfies the required conditions. The definition of these conditions is based on a sequence of control function $\mathrm{L}_{t,T}^{0}$ and $\mathrm{L}_{t,T}^{1}$, which have backward recursive definitions for $t$ from $T$ to $1$.
@@
In Section 10.2, we summarize the reduction introduced in (lyu2022composition, ), with the control functions $\mathrm{L}_{t,T}^{b}$ replaced by $\mathrm{L}_{t}^{b}$. We also modify this reduction by adding a new condition to the original three that the functions $\phi_{t}$ must satisfy. This additional condition is used to limit the number of interactions between $\mathcal{P}$ and $\mathrm{IRR}_{\epsilon,\delta}(b)$. Finally, in Section 10.3, we provide a new construction of $\phi_{t}$ and show that it satisfies all four conditions.
@@
Lyu (lyu2022composition, ) introduces the following control function:
@@
Lyu (lyu2022composition, ) shows the following two claims about $\mathrm{Lower}_{t,T}^{b}$:
Claim 10.5 ((lyu2022composition, )).
For each $b\in\{0,1\}$, $T\in\mathbb{N}$, and $q_{1}\in Q$, we have $\mathrm{Lower}_{1,T}^{b}(q_{1})\leq\delta$.
Claim 10.6 ((lyu2022composition, )).
For each $b\in\{0,1\}$, $T\in\mathbb{N}$, $t\in[T]$, [^13] In the original claim in (lyu2022composition, ), $t$ is assumed to be at most T1T-1, which is a typo. and transcript $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{t-1}\times Q$, we have
@@
For each $b\in\{0,1\}$ and $T\in\mathbb{N}$, we have $\mathrm{L}_{0,T}^{b}()\leq\delta$.
In some part of the proof, Lyu (lyu2022composition, ) wrongly refers to Claim 10.6, while the following lemma is required. For completeness, we prove this lemma and use it in the next section.
@@
This section introduces our main new technical contribution. To remove the assumption of having an upper bound $T$ on the number of queries, we introduce $\mathrm{L}_{t}^{b}$ as the limit of $\mathrm{L}_{t,T}^{b}$ and show that it satisfies all properties needed by the proof of (lyu2022composition, ). Thus, differently from (lyu2022composition, ), we will use $\mathrm{L}_{t}^{0}$ and $\mathrm{L}_{t}^{1}$ to define the function $\phi_{t}$.
@@
To ensure $\mathrm{L}_{t}^{b}$ is well-defined, we need to show that the limit of $\mathrm{L}_{t,T}^{b}$ exists. We will upper bound the values of $\mathrm{L}_{t,T}^{b}$ by $1$ and show that this function is non-decreasing in $T$. As a result, for each $t\in\mathbb{N}\cup\{0\}$ and $(q_{1},\dots,a_{t})\in(Q\times A)^{t}$, the limit $\lim_{T\geq t,T\to\infty}\mathrm{L}_{t,T}^{b}(q_{1},\dots,a_{t})$ exists. We need the following lemma throughout the section:
@@
$=\sum\limits_{a_{T}\in A}\sup_{q_{T+1}\in Q}\max\{0,\sum\limits_{a_{T+1}\in A}\mu_{T+1}^{b}(q_{1},\dots,a_{T+1})-e^{\epsilon}\mu_{T+1}^{1-b}(q_{1},\dots,a_{T+1})\}$
@@
In this section, we present the properties of $\mathrm{L}_{t}^{b}$, required in the next sections. The definition of $\mathrm{L}_{t}^{b}$ combined with Corollary 10.7, Lemma 10.8, and Lemma 10.12 gives the following corollaries, respectively:
@@
For each $b\in\{0,1\}$, we have $\mathrm{L}_{0}^{b}()\leq\delta$.
@@
To define $\mathcal{P}$ and simulate $\mathcal{M}$, Lyu (lyu2022composition, ) constructs four families of probability mass functions (PMFs), corresponding to each outcome of $\mathrm{RR}_{\epsilon,\delta}$. Intuitively, each family describes an IM and consists of PMFs over the answer set $A$ for every $t\in[T]$ and every history $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{t}\times Q$. These PMFs represent the probability of returning each answer in response to $q_{t}$, conditioned on $q_{1},\dots,a_{t-1}$ being the history of queries and answers. Similarly, we define $\mathcal{P}$ based on four families of PMFs: $\mathcal{F}_{(\bot,0)}$, $\mathcal{F}_{(\bot,1)}$, $\mathcal{F}_{(\top,0)}$, and $\mathcal{F}_{(\top,1)}$, corresponding to the four possible output sequences of $\mathrm{IRR}_{\epsilon,\delta}$. We note that in our construction, each family includes a PMF over $A$ for every $t\in\mathbb{N}$ and every history $(q_{1},\dots,a_{t-1},q_{t})$. Moreover, the definitions of these PMFs are not the same as the ones in (lyu2022composition, ).
@@
In (lyu2022composition, ), the post-processing mechanism initially interacts with $\mathrm{RR}_{\epsilon,\delta}$ and stores its output $r=(\tau,c)\in\{\top,\bot\}\times\{0,1\}$. It then answers the adversary’s queries by itself using the family of PMFs corresponding to $r$. Our high-level idea is to construct the families $\mathcal{F}_{(\bot,0)}$, $\mathcal{F}_{(\bot,1)}$, $\mathcal{F}_{(\top,0)}$, and $\mathcal{F}_{(\top,1)}$ in such a way that, under certain conditions, knowledge of $\tau$ alone suffices to choose the appropriate PMF. Recall that the IPM $\mathcal{P}$ interacts with $\mathrm{IRR}_{\epsilon,\delta}$ instead of $\mathrm{RR}_{\epsilon,\delta}$. We design $\mathcal{P}$ to initially interact with $\mathrm{IRR}_{\epsilon,\delta}$ and store its response. $\mathcal{P}$ then interacts with $\mathrm{IRR}_{\epsilon,\delta}$ for the second time when it cannot choose a PMF using only the first response.
@@
To generate a response $a$ to query $q$, $\mathcal{P}$ proceeds as follows: If $q$ is the first query (i.e., $r=()$), then $\mathcal{P}$ initially interacts with $\mathrm{IRR}_{\epsilon,\delta}$ and appends its response to $r$. After this, whether $q$ is the first query or not, the first time the following condition does not hold, $\mathcal{P}$ interacts with $\mathrm{IRR}_{\epsilon,\delta}$ for the second time and appends its answer to $r$: the PMFs corresponding to $(h,q)$ in $\mathcal{F}_{(\top,0)}$ and $\mathcal{F}_{(\top,1)}$ are identical, and the PMFs corresponding to $(h,q)$ in $\mathcal{F}_{(\bot,0)}$ and $\mathcal{F}_{(\bot,1)}$ are identical. After that, $\mathcal{P}$ never interacts with $\mathrm{IRR}_{\epsilon,\delta}$ again.
@@
Using the control function $\mathrm{Lower}_{t,T}$, Lyu (lyu2022composition, ) introduces three conditions and constructs a sequence of functions $\phi_{t}:(Q\times A)^{t}\to[0,1]^{2}$ for $t\in\{1,\dots,T\}$ that satisfy these conditions. The families of PMFs are defined then based on these functions. We remove the dependence on the bound $T$ by replacing $\mathrm{Lower}_{t,T}$ with $\mathrm{L}_{t}$ in the conditions, and extend the construction of $\phi_{t}$ to all $t\in\mathbb{N}$. Furthermore, we introduce an additional condition to those given in (lyu2022composition, ), which we will later use to bound the number of times $\mathcal{P}$ interacts with $\mathrm{IRR}_{\epsilon,\delta}$.
@@
Similar to the previous section, $\mathcal{F}_{(\top,0)}$ and $\mathcal{F}_{(\top,1)}$ are defined based on a sequence of functions $\psi_{t}:(Q\times A)^{t}\to[0,1]^{2}$ for each $t\in\mathbb{N}\cup\{0\}$. For $t=0$, $\psi_{t}()=(1,1)$. For $t\in\mathbb{N}$, we define $\psi_{t}$ as follows:
@@ 10.3. Construction of $\phi_{t}$
For each $q_{1},\dots,a_{t-1},q_{t}$, Lyu (lyu2022composition, ) initializes the values of $\phi_{t}(q_{1},\dots,a_{t-1},q_{t},\cdot)$ with the ones from $\mathrm{L}_{t,T}(q_{1},\dots,a_{t-1},q_{t},\cdot)$. (Removing the upper bound assumption, they are initialized by the values of $\mathrm{L}_{t}(q_{1},\dots,a_{t-1},q_{t},\cdot)$.) Then, they iterate over all possible answers $a_{t}\in A$ in a fixed order and gradually increase the value of $\phi_{t}(q_{1},\dots,a_{t})$ to satisfy conditions (II) and (III), while ensuring that condition (I) remains valid. This iterative procedure requires $A$ to be finite; otherwise, the construction algorithm does not terminate. Moreover, the fixed iteration order can lead to an imbalance in which earlier values grow significantly larger than those that come later in the order. In our new construction, we control and balance this increase.
@@
Lemma 10.25 (Transfinite Recursion (suppes2012axiomatic, )).
@@ 10.3.4. Satisfying Conditions (I) to (IV)
Our objective was to construct $\phi_{t}^{b}$ so that it fulfills conditions (I) to (IV). Combination of Constraints ($\phi$.1) and ($\xi$.1) gives condition (I). Moreover, Constraint ($\phi$.3) directly implies condition (II). The proof of the following claim is identical to a proof in (lyu2022composition, ):
@@
which is a contradiction when $\epsilon>0$, implying that condition (III) holds. ∎
@@
for all $(a_{1},\dots,a_{t})\in A^{t}$. Moreover, by Lemma 10.10 and $\star$10.3.4, $\mu_{t-1}^{0}(q_{1},\dots,a_{t-1})=\mu_{t-1}^{1}(q_{1},\dots,a_{t-1})$ for all $(a_{1},\dots,a_{t})\in A^{t}$. Thus, by Remark 10.21 and condition (IV),
@@
Therefore, by equations ($\star$10.3.4), ($\star\star$10.3.4) and ($\star\star\star$10.3.4), Constraints ($\phi$.1), ($\phi$.2), and ($\phi$.3) are symmetric for $x_{0}$ and $x_{1}$. Hence, there exists a maximal feasible solution $(z^{*},z^{*})$ satisfying all the constraints. By definition of $\phi_{t}(q_{1},\dots,a_{t})$, we have $\phi_{t}(q_{1},\dots,a_{t})=(z^{*},z^{*})$. Also, by Constraint ($\phi$.1), $z^{*}$ is at least $\min\{\mu_{t}^{0}(q_{1},\dots,a_{t}),\mathrm{L}_{t}^{0}(q_{1},\dots,a_{t})+\mathrm{L}_{t}^{1}(q_{1},\dots,a_{t})\}$.